Given the following data:

Solution

Conc. (M)

Volume (mL)

Tinitial (ºC)

Tfinal (ºC)

NaOH

0.750

50.00

19.84

23.18

HCl

0.500

50.00

19.84

Calculate ∆Hº for the reaction:

            HCl(aq) + NaOH(aq) → H2O(l) + NaCl(aq)

You may assume all solutions to have the same density and specific heat as water (1.00 g/mL and 4.184 J/gºC)

Answer in kJ/mol
Answer #1

HCl(aq) + NaOH(aq) → H2O(l) + NaCl(aq)

total volume = 50 +50 = 100ml

mass of solution = volume * density

                         = 100*1   = 100g

q   = mc\DeltaT

     = 100*4.184*(23.18-19.84)   = 1397.45J

∆Hº   = 1397.45J

no of moles of HCl = molarity * volume in L

                                 = 0.5*0.05   = 0.025 moles

no of moles of NaOH   = molarity * volume in L

                                    = 0.75*0.05 = 0.0375mole

HCl is limiting reactant

the reaction enthalpy per mole   = 1397.45/0.025   = 55898J/mole = 55.898Kj/mole